# Enhancing Reasoning for Diffusion LLMs via Distribution Matching Policy Optimization

### v1

@@ Abstract
Diffusion large language models (dLLMs) are promising alternatives to autoregressive large language models (AR-LLMs), as they potentially allow higher inference throughput. Reinforcement learning (RL) is a crucial component for dLLMs to achieve comparable performance with AR-LLMs on important tasks, such as reasoning. However, RL algorithms that are well-suited for dLLMs’ unique characteristics have yet to be developed. This paper proposes Distribution Matching Policy Optimization (DMPO), a principled and theoretically grounded RL fine-tuning method specifically designed to enhance the reasoning capabilities of dLLMs by matching the dLLM policy distribution to the optimal, reward-tilted one through cross-entropy optimization. We identify a key challenge in the implementation with a small training batch size and propose several effective solutions through a novel weight baseline subtraction technique. DMPO exhibits superior performance on multiple reasoning benchmarks without supervised fine-tuning, with an accuracy improvement of up to $42.9\%$ over previously SOTA baselines and $55.8\%$ over the base model, underscoring the effectiveness of the distribution matching framework. Our code is available at https://github.com/yuchen-zhu-zyc/DMPO.
@@ 1 Introduction
With the aim of addressing such issues, diffusion large language models (dLLMs) have been investigated as an alternative to the AR models. Unlike their counterparts, dLLMs iteratively refine a sequence from a masked state, allowing for any-order generation, and have shown promising performance in text generation tasks. dLLMs, such as LLaDA (Nie et al., 2025b) and Dream (Ye et al., 2025), have demonstrated competitive performances on many tasks compared to similar-size AR baselines. Recently, commercial models such as Mercury (Inception Labs et al., 2025) and Gemini Diffusion (DeepMind, ) have demonstrated the capability to achieve a magnitude higher inference throughput without sacrificing generation quality, suggesting dLLM as a promising future direction for language modeling. However, one question that remains largely unanswered is how to transfer the success of RL on LLM to dLLM, thereby further scaling up the model’s skills.
@@
Designing RL algorithms for dLLMs faces two major challenges. Due to the bidirectional nature of dLLMs, estimating the log probability of the generated sequences is more expensive than for AR models, making it less favorable to naively adapt LLM post-training algorithms like GRPO to dLLMs, as they heavily rely on such estimation. The GRPO-style algorithms also do not leverage dLLM’s unique characteristic of having a forward noising process, as they are backward-only algorithms when using generated rollouts. Moreover, existing RL frameworks for enhancing LLM reasoning capabilities overly focus on reward maximization (Guo et al., 2025a; Liu et al., 2025c; Zheng et al., 2025a). By targeting only the reward mode, these approaches do not properly utilize dLLM’s potential in generating more diverse responses than LLMs due to the random-order nature (Gong et al., 2025).
@@
To jointly address these challenges, we propose Distribution Matching Policy Optimization (DMPO), a principled and efficient RL fine-tuning method specifically designed for dLLMs. DMPO is designed based on a novel framework theoretically grounded in stochastic optimal control (SOC), which shifts away from the conventional reward maximization paradigm and targets a new goal of matching the entire reward-tilted policy distribution. This enables the model to explore diverse, high-quality reasoning paths and responses during training, addressing concerns about over-focusing on absolute reward values and modes. Moreover, DMPO training leverages importance sampling and a novel weighted denoising cross-entropy (WDCE) loss, which enjoys the key advantage of operating in an off-policy manner, allowing the use of replay buffers for improved sample efficiency. More importantly, WDCE is a forward-only objective that relies solely on the obtained clean samples and the inexpensive, forward-noising process unique to diffusion LLMs. DMPO largely discards the dependence on rollout trajectories, enabling it to potentially enjoy more speed-up than other dLLM RL algorithms when employed with fast inference techniques.
@@
Contributions. The core contributions of this paper are summarized as follows: (I) We propose a novel RL learning framework for dLLMs that targets distribution matching rather than reward maximization (Sec.˜3.1). (II) We propose Distribution Matching Policy Optimization (DMPO), a principled, theoretically-grounded fine-tuning strategy for enhancing dLLM’s reasoning capabilities, supported by importance sampling and weighted denoising cross-entropy (Sec.˜3.2). (III) We identify a special challenge that occurred for WDCE due to the use of a limited training batch size, and propose two novel techniques to address it: weight baseline subtraction (Sec.˜3.3) and weighted direct discriminative optimization (Sec.˜3.4). (IV) DMPO exhibits superior performances on multiple reasoning benchmarks without SFT, with an accuracy improvement up to $42.9\%$ over previously SOTA baselines and $55.8\%$ over the base model, being top-performing across bi-directional dLLMs (Sec.˜4).
Figure 1: Performances on reasoning benchmarks evaluated with generation length $256$. DMPO constantly achieves the best performances across bidirectional dLLM, outperforming d1.
@@ 2.1 Masked Diffusion Models for Language Modeling
$\Pr_{{\bm{X}}\sim p_{\mathrm{data}}}(X_{d}=u|{\bm{X}}_{\mathrm{UM}}={\bm{x}}_{\mathrm{UM}}),\quad\text{where}~{\bm{x}}_{\mathrm{UM}}=(x_{d}:x_{d}\neq\mathsf{M}).$
By definition, we assume each row of ${\bm{\pi}}_{\theta}({\bm{x}})$ is a valid probability vector. The probability of a unmasked sequence ${\bm{x}}\in\mathcal{V}^{D}$ under the MDM ${\bm{\pi}}_{\theta}$ is defined through random-order autoregressive (AR) generation: choosing a uniformly random order of the $D$ positions, and autoregressively sampling each position conditional on the previously sampled ones. Formally,
$p_{\theta}({\bm{x}})=\operatorname{\mathbb{E}}_{\bm{\sigma}}p_{\theta}({\bm{x}};{\bm{\sigma}}),\quad\text{where}~{\bm{\sigma}}\sim\operatorname{Unif}(S_{D})~\text{and}~p_{\theta}({\bm{x}};{\bm{\sigma}})=\prod_{d=1}^{D}{\bm{\pi}}_{\theta}(x_{\sigma_{d}}|{\bm{x}}_{{\bm{\sigma}}_{<d}}),$ (1)
where $S_{D}$ is the set of all permutations of $\{1,...,D\}$; ${\bm{\pi}}_{\theta}(x_{\sigma_{d}}|{\bm{x}}_{{\bm{\sigma}}_{<d}})$ means input $\bm{x}$} with all positions except ${\bm{\sigma}}_{<d}=\{\sigma_{1},...,\sigma_{d-1}\}$ masked into the MDM and take the output at position $(\sigma_{d},x_{\sigma_{d}})$.
The standard way for training a masked discrete diffusion model given i.i.d. samples from $p_{\mathrm{data}}$ is minimizing the denoising cross-entropy (DCE) loss $\operatorname{\mathbb{E}}_{p_{\mathrm{data}}({\bm{x}})}\mathcal{L}_{\theta}({\bm{x}})$, which involves the following definition of the (negative) evidence lower bound (ELBO) $\mathcal{L}_{\theta}$:
$-\log p_{\theta}({\bm{x}})$ $=-\log\operatorname{\mathbb{E}}_{{\bm{\sigma}}}p_{\theta}({\bm{x}};{\bm{\sigma}})\leq-\operatorname{\mathbb{E}}_{{\bm{\sigma}}}\log p_{\theta}({\bm{x}};{\bm{\sigma}})\qquad(\textit{Jensen's inequality})$
$=\operatorname{\mathbb{E}}_{m\sim\operatorname{Unif}\{1,...,|{\bm{x}}|\}}\bigg[\frac{|{\bm{x}}|}{m}\operatorname{\mathbb{E}}_{\mu_{m}({\widetilde{\bm{x}}}|{\bm{x}})}\sum_{d:\widetilde{x}_{d}=\mathsf{M}}-\log{\bm{\pi}}_{\theta}({\widetilde{\bm{x}}})_{d,x_{d}}\bigg]=:\mathcal{L}_{\theta}({\bm{x}}),$ (2)
@@
When applying to text data, the MDM is also referred to as the diffusion large language model (dLLM) (Nie et al., 2025b; Ye et al., 2025; Inception Labs et al., 2025; Song et al., 2025). For the purpose of reasoning, we typically write ${\bm{x}}=({\bm{q}},{\bm{o}})$, where $\bm{q}$} is the prompt (or query, which is always assumed to contain no mask state) and $\bm{o}$} is the response (or output). We use $\pi_{\theta}({\bm{o}}|{\bm{q}})\in\mathbb{R}^{|{\bm{o}}|\times V}$ to denote the policy model output of the dLLM given a prompt $\bm{q}$} and a partially masked response $\bm{o}$}. The conditional sequence probability of a clean model $\bm{o}$} given a prompt $\bm{q}$}, denoted as $p_{\theta}({\bm{o}}|{\bm{q}})$, is similarly defined through ˜1, where we now use notations $p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ and ${\bm{\pi}}_{\theta}(o_{\sigma_{d}}|{\bm{q}},{\bm{o}}_{{\bm{\sigma}}_{<d}})$ to emphasize the dependence on the prompt $\bm{q}$}. The negative ELBO will be written as $\mathcal{L}_{\theta}({\bm{o}}|{\bm{q}})$.
@@ 2.2 Reinforcement Learning for Enhancing Reasoning
We first present the Group Relative Policy Optimization (GRPO, Shao et al. (2024)) method for LLMs, which is the basis of most of the existing RL methods for dLLMs. Given a pretrained LLM with policy $\pi_{\mathrm{ref}}$ that samples from the distribution $p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})=\prod_{d=1}^{|{\bm{o}}|}\pi_{\mathrm{ref}}(o_{d}|{\bm{q}},{\bm{o}}_{<d})$, a reward function $r:({\bm{q}},{\bm{o}})\mapsto\mathbb{R}$, a set of prompts $\mathcal{D}$, and a regularization parameter $\alpha\geq 0$, each step of the GRPO aims to solve the following problem: sample ${\bm{q}}\sim\mathcal{D}$, ${\bm{o}}^{(1:G)}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}p_{\theta_{\mathrm{old}}}({\bm{o}}|{\bm{q}})$, and maximize
$\operatorname{\mathbb{E}}\bigg\{\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|{\bm{o}}^{(i)}|}\sum_{d=1}^{|{\bm{o}}^{(i)}|}\left[\min\left(\rho_{d}^{(i)}A_{i},\operatorname{clip}(\rho_{d}^{(i)})_{1\pm\epsilon}A_{i}\right)-\alpha\operatorname{KL}(p_{\theta}({\bm{o}}^{(i)}|{\bm{q}})\|p_{\mathrm{ref}}({\bm{o}}^{(i)}|{\bm{q}}))\right]\bigg\},$ (3)
where the advantages [^1] As suggested by Liu et al. (2025c), we list here the version without normalization by standard deviation. are $A_{i}=r({\bm{q}},{\bm{o}}^{(i)})-\mathrm{mean}(r({\bm{q}},{\bm{o}}^{(1:G)}))$, the per-token probability ratios are $\rho_{d}^{(i)}=\frac{\pi_{\theta}(o_{d}^{(i)}|{\bm{q}},{\bm{o}}_{<d}^{(i)})}{\pi_{\theta_{\mathrm{old}}}(o_{d}^{(i)}|{\bm{q}},{\bm{o}}_{<d}^{(i)})}$, and the KL regularization term is estimated similarly by the per-token probability ratios between $\pi_{\theta}$ and $\pi_{\mathrm{ref}}$. $\epsilon$ is a clipping threshold that prevents overly large policy updates.
While ˜3 works well for LLMs, it is not directly applicable to dLLMs due to mismatch between the dLLM policy (model output) ${\bm{\pi}}_{\theta}({\bm{o}}|{\bm{q}})$ and the sequence likelihood $p_{\theta}({\bm{o}}|{\bm{q}})$: unlike in LLMs where these two quantities are easily connected through the chain rule, it is generally non-trivial to compute the per-token probability given the dLLM model output, and only ELBO ˜2 is available as a surrogate. To tackle this issue, diffu-GRPO (Zhao et al., 2025a) proposed the following strategy: first, one fully masks all response positions except $d$, and partially masks the prompt $\bm{q}$}. By feeding this sequence into the model, one obtains the approximate probability $p_{\theta}(o_{d}|{\bm{q}})$. Second, the sequence probability $p_{\theta}({\bm{o}}|{\bm{q}})$ is approximated by mean-field decomposition: $p_{\theta}({\bm{o}}|{\bm{q}})\approx\prod_{d=1}^{|{\bm{o}}|}p_{\theta}(o_{d}|{\bm{q}})$. Such approximations do not capture the correlation between different positions in the response, which produces imprecision. A similar technique is employed in coupled-GRPO for code generation tasks (Gong et al., 2025).

### v2

@@ Abstract
Diffusion large language models (dLLMs) are promising alternatives to autoregressive large language models (AR-LLMs), as they potentially allow higher inference throughput. Reinforcement learning (RL) is a crucial component for dLLMs to achieve comparable performance with AR-LLMs on important tasks, such as reasoning. However, RL algorithms that are well-suited for dLLMs’ unique characteristics have yet to be developed. This paper proposes Distribution Matching Policy Optimization (DMPO), a principled and theoretically grounded RL fine-tuning method specifically designed to enhance the reasoning capabilities of dLLMs by matching the dLLM policy distribution to the optimal, reward-tilted one through cross-entropy optimization. We identify a key challenge in the implementation with a small training batch size and propose several effective solutions through a novel weight baseline subtraction technique. DMPO exhibits superior performance on multiple reasoning benchmarks without supervised fine-tuning, with an accuracy improvement of up to $54.3\%$ over previously SOTA baselines and $66.41\%$ over the base model, underscoring the effectiveness of the distribution matching framework. Our code is available at https://github.com/yuchen-zhu-zyc/DMPO.
Machine Learning, ICML
Figure 1: Performance on reasoning benchmarks evaluated with generation length $256$. DMPO consistently achieves the best performance across dLLMs, outperforming d1 and cGRPO.
@@ 1 Introduction
To address such issues, diffusion large language models (dLLMs) have been investigated as an alternative to the AR models. Unlike AR-LLMs, dLLMs iteratively refine a sequence from masks, allowing for any-order generation, and have shown promising performance on text modeling tasks. dLLMs such as LLaDA (Nie et al., 2025b) and Dream (Ye et al., 2025), have demonstrated competitive performances on many tasks compared to similar-size AR baselines. Recently, commercial models such as Mercury (Inception Labs et al., 2025) and Gemini Diffusion (DeepMind, ) have demonstrated the capability to achieve a magnitude higher inference throughput without sacrificing generation quality, suggesting dLLM as a promising direction for language modeling. However, one question that remains largely unanswered is how to transfer the success of RL on LLMs to dLLMs, thereby further scaling up their capability.
@@
Designing RL algorithms for dLLMs faces two major challenges. Due to the bidirectional nature of dLLMs, estimating the log probability of the generated sequences is more expensive than for AR models, making it less favorable to naively adapt LLM post-training algorithms like GRPO to dLLMs, as they heavily rely on such estimation. The GRPO-style algorithms also do not leverage dLLM’s unique characteristic of having a forward noising process, as they are backward-only algorithms when using generated rollouts. Moreover, existing RL frameworks for enhancing LLM reasoning capabilities overly focus on reward maximization (Guo et al., 2025a; Liu et al., 2025c; Zheng et al., 2025a). By targeting only the reward mode, these approaches do not properly utilize dLLMs’ potential in generating diverse responses due to the random-order nature (Gong et al., 2025).
@@
To jointly address these challenges, we propose Distribution Matching Policy Optimization (DMPO), a principled and efficient RL fine-tuning method specifically designed for dLLMs. DMPO is designed based on a novel framework grounded in stochastic optimal control (SOC), which shifts away from the conventional reward maximization paradigm and targets a new goal of matching the entire reward-tilted policy distribution. This enables the model to explore diverse, high-quality reasoning paths and responses during training, addressing concerns about over-focusing on absolute reward values and modes. In addition, DMPO training leverages importance sampling and a novel weighted denoising cross-entropy (WDCE) loss, which enjoys the key advantage of operating in an off-policy manner, allowing the use of replay buffers for improved sample efficiency. More importantly, WDCE is a forward-only objective that relies solely on the obtained clean samples and the inexpensive, forward-noising process unique to dLLMs. DMPO largely discards the dependence on rollout trajectories, enabling it to potentially enjoy more speed-up than other dLLM RL algorithms when employed with fast inference techniques.
Contributions
@@
(I) We propose a novel RL learning framework for dLLMs that targets distribution matching rather than reward maximization (Sec. 3.1). (II) We propose Distribution Matching Policy Optimization (DMPO), a principled, theoretically-grounded fine-tuning strategy for enhancing dLLM’s reasoning capabilities, supported by importance sampling and weighted denoising cross-entropy (Sec. 3.2). (III) We identify a special challenge that occurred for WDCE due to the use of a limited training batch size, and propose two novel techniques to address it: weight baseline subtraction (Sec. 3.3) and weighted direct discriminative optimization (Sec. 3.4). (IV) DMPO exhibits superior performances on multiple reasoning benchmarks without supervised fine-tuning (SFT), with an accuracy improvement up to $54.3\%$ over previously SOTA baselines and $66.41\%$ over the base model, being top-performing across bi-directional dLLMs (Sec. 4).
@@ 2.1 Masked Diffusion Models for Language Modeling
$\Pr\nolimits_{{\bm{X}}\sim p_{\mathrm{data}}}(X_{d}=u|{\bm{X}}_{\mathrm{UM}}={\bm{x}}_{\mathrm{UM}}),$
where ${\bm{x}}_{\mathrm{UM}}=(x_{d}:x_{d}\neq\mathsf{M})$. By definition, we assume each row of ${\bm{\pi}}_{\theta}({\bm{x}})$ is a valid probability vector. The probability of a unmasked sequence ${\bm{x}}\in\mathcal{V}^{D}$ under the MDM ${\bm{\pi}}_{\theta}$ is defined through random-order autoregressive (ROAR) generation: choosing a uniformly random order of the $D$ positions, and autoregressively sampling each position conditional on the previously sampled ones. Formally, let $\bm{\sigma}$} be a uniformly random permutation of $\{1,...,D\}$, then
$p_{\theta}({\bm{x}})=\operatorname{\mathbb{E}}_{{\bm{\sigma}}}p_{\theta}({\bm{x}};{\bm{\sigma}}),~p_{\theta}({\bm{x}};{\bm{\sigma}})=\prod_{d=1}^{D}{\bm{\pi}}_{\theta}(x_{\sigma_{d}}|{\bm{x}}_{{\bm{\sigma}}_{<d}}).$ (1)
Here, ${\bm{\pi}}_{\theta}(x_{\sigma_{d}}|{\bm{x}}_{{\bm{\sigma}}_{<d}})$ means input $\bm{x}$} with all positions except ${\bm{\sigma}}_{<d}=\{\sigma_{1},...,\sigma_{d-1}\}$ masked into the MDM and take the output at position $(\sigma_{d},x_{\sigma_{d}})$.
The standard way to train an MDM given i.i.d. samples from $p_{\mathrm{data}}$ is to minimize the denoising cross-entropy (DCE) loss $\operatorname{\mathbb{E}}_{p_{\mathrm{data}}({\bm{x}})}\mathcal{L}_{\theta}({\bm{x}})$, which involves the following definition of the (negative) evidence lower bound (ELBO) $\mathcal{L}_{\theta}$:
$-\log p_{\theta}({\bm{x}})=-\log\operatorname{\mathbb{E}}_{{\bm{\sigma}}}p_{\theta}({\bm{x}};{\bm{\sigma}})\leq-\operatorname{\mathbb{E}}_{{\bm{\sigma}}}\log p_{\theta}({\bm{x}};{\bm{\sigma}})$
$=\operatorname{\mathbb{E}}_{m\sim\operatorname{Unif}\{1,...,|{\bm{x}}|\}}\Big[\frac{|{\bm{x}}|}{m}\operatorname{\mathbb{E}}_{\mu_{m}({\widetilde{\bm{x}}}|{\bm{x}})}\sum_{d:\widetilde{x}_{d}=\mathsf{M}}-\log{\bm{\pi}}_{\theta}({\widetilde{\bm{x}}})_{d,x_{d}}\Big]=:\mathcal{L}_{\theta}({\bm{x}}),$ (2)
@@
When applying to text data, the MDM is also referred to as the diffusion large language model (dLLM) (Nie et al., 2025b; Ye et al., 2025; Inception Labs et al., 2025; Song et al., 2025). For the purpose of reasoning, we typically write ${\bm{x}}=({\bm{q}},{\bm{o}})$, where $\bm{q}$} is the prompt (or query, which is always assumed to contain no mask state) and $\bm{o}$} is the response (or output). We use $\pi_{\theta}({\bm{o}}|{\bm{q}})\in\mathbb{R}^{|{\bm{o}}|\times V}$ to denote the policy model output of the dLLM given a prompt $\bm{q}$} and a partially masked response $\bm{o}$}. The conditional sequence probability of a clean model $\bm{o}$} given a prompt $\bm{q}$}, denoted as $p_{\theta}({\bm{o}}|{\bm{q}})$, is similarly defined through 1, where we now use notations $p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ and ${\bm{\pi}}_{\theta}(o_{\sigma_{d}}|{\bm{q}},{\bm{o}}_{{\bm{\sigma}}_{<d}})$ to emphasize the dependence on the prompt $\bm{q}$}. The negative ELBO will be written as $\mathcal{L}_{\theta}({\bm{o}}|{\bm{q}})$.
@@ 2.2 Reinforcement Learning for Enhancing Reasoning
We first present the Group Relative Policy Optimization (GRPO, Shao et al. (2024)) method for LLMs, which is the basis of most of the existing RL methods for dLLMs. Given a pretrained LLM with policy $\pi_{\mathrm{ref}}$ that samples from the distribution $p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})=\prod_{d=1}^{|{\bm{o}}|}\pi_{\mathrm{ref}}(o_{d}|{\bm{q}},{\bm{o}}_{<d})$, a reward function $r:({\bm{q}},{\bm{o}})\mapsto\mathbb{R}$, a set of prompts $\mathcal{D}$, and a regularization parameter $\alpha\geq 0$, each step of the GRPO aims to solve the following problem: sample prompt ${\bm{q}}\sim\mathcal{D}$ and rollouts ${\bm{o}}^{(1:G)}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}p_{\theta_{\mathrm{old}}}({\bm{o}}|{\bm{q}})$, then, maximize
$\operatorname{\mathbb{E}}\Big\{\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|{\bm{o}}^{(i)}|}\sum_{d=1}^{|{\bm{o}}^{(i)}|}\Big[\min\Big(\rho_{d}^{(i)}A_{i},\operatorname{clip}(\rho_{d}^{(i)})_{1\pm\epsilon}A_{i}\Big)-\alpha\operatorname{KL}(p_{\theta}({\bm{o}}^{(i)}|{\bm{q}})\|p_{\mathrm{ref}}({\bm{o}}^{(i)}|{\bm{q}}))\Big]\Big\},$ (3)
where the advantages [^1] As suggested by Liu et al. (2025c), we list here the version without normalization by standard deviation. are $A_{i}=r({\bm{q}},{\bm{o}}^{(i)})-\mathrm{mean}(r({\bm{q}},{\bm{o}}^{(1:G)}))$, the per-token probability ratios are $\rho_{d}^{(i)}=\frac{\pi_{\theta}(o_{d}^{(i)}|{\bm{q}},{\bm{o}}_{<d}^{(i)})}{\pi_{\theta_{\mathrm{old}}}(o_{d}^{(i)}|{\bm{q}},{\bm{o}}_{<d}^{(i)})}$, and the KL regularization term is estimated similarly by the per-token probability ratios between $\pi_{\theta}$ and $\pi_{\mathrm{ref}}$. The clipping threshold $\epsilon$ prevents overly large policy updates.
While 3 works well for LLMs, it is not directly applicable to dLLMs due to mismatch between the dLLM policy (model output) ${\bm{\pi}}_{\theta}({\bm{o}}|{\bm{q}})$ and the sequence likelihood $p_{\theta}({\bm{o}}|{\bm{q}})$: unlike in LLMs where these two quantities are easily connected through the chain rule, it is generally non-trivial to compute the per-token probability given the dLLM model output, and only ELBO 2 is available as a surrogate. To tackle this issue, diffu-GRPO (Zhao et al., 2025a) proposed to fully mask all response positions and partially masks the prompt $\bm{q}$}, and feed this sequence into the model to obtain the approximate probability $p_{\theta}(o_{d}|{\bm{q}})$. Next, the sequence probability $p_{\theta}({\bm{o}}|{\bm{q}})$ is approximated by mean-field decomposition: $p_{\theta}({\bm{o}}|{\bm{q}})\approx\prod_{d=1}^{|{\bm{o}}|}p_{\theta}(o_{d}|{\bm{q}})$. Such approximations do not capture the correlation between different positions in the response, which produces imprecision. A similar technique is employed in coupled-GRPO (cGRPO) for code generation tasks in Gong et al. (2025).